# Neural Collapse-Inspired Multi-Label Federated Learning under Label-Distribution Skew

### v2

@@ Introduction
$\mathbf{h}_{ic}\;=\;\operatorname{MultiHeadAttn}(\mathbf{m}_{c},\;\mathbf{f}_{i}),\quad c\in\mathcal{C}$ (3)
and stacks them into
@@
Sharing the fixed query matrix $\mathbf{M}$ across clients mitigates aggregation conflicts and encourages globally aligned, disentangled features.
Neural Collapse-Inspired Feature Alignment
After disentangling image-level features into class-wise representations with LADM, each sample $i\in[N_{k}]$ yields a feature $\mathbf{h}_{ic}\in\mathbb{R}^{d}$ for every class $c\in\mathcal{C}$. To avoid client-specific classifier drift caused by label imbalance or missing classes, we fix the classifier to the global simplex ETF used as the LADM query matrix, $\mathbf{M}=[\mathbf{m}_{c}]_{c=1}^{C}\in\mathbb{R}^{d\times C}$. Given $\mathbf{h}_{ic}$ and its corresponding prototype $\mathbf{m}_{c}$, we compute a class logit and pass it through the sigmoid to obtain a binary prediction:
@@
and train with a binary cross-entropy (BCE) objective
@@
Here, $y_{ic}\in\{0,1\}$ is the ground-truth multi-label indicator, $\sigma(\cdot)$ denotes the sigmoid. This fixed classifier improves class separability and reduces divergence between local and global label distributions.
Reducing Noise and Improving Clustering
Each sample produces $C$ class-wise features $\{\mathbf{h}_{ic}\}_{c=1}^{C}$. Let the positive and negative class sets for sample $i$ be $\mathcal{C}_{i}^{+}=\{\,c\mid y_{ic}=1\,\}$ and $\mathcal{C}_{i}^{-}=\{\,c\mid y_{ic}=0\,\}$, respectively. We introduce two regularizers: one to suppress noise from negative features and another to encourage compact clustering of positive features.
Negative Feature Rejection Loss.
While $\mathcal{L}_{\text{BCE}}$ discourages alignment of a negative feature $\mathbf{h}_{ic}$ with its own prototype $\mathbf{m}_{c}$ when $y_{ic}=0$, it does not prevent $\mathbf{h}_{ic}$ from spuriously aligning with other prototypes. We therefore penalize high similarity between every negative feature and all non-self prototypes:
@@
$\mathcal{L}_{\text{Neg}}=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{-}|}\sum_{c\in\mathcal{C}_{i}^{-}}\frac{1}{C-1}\sum_{\begin{subarray}{c}r=1\\ r\neq c\end{subarray}}^{C}\mathbb{I}\!\big(\hat{s}_{icr}>\tau\big)\,\log\!\big(1-\hat{s}_{icr}\big).$ (8)
The indicator $\mathbb{I}(\cdot)$ filters out low-similarity pairs so that only confident negatives contribute. In our experiments, we set $\tau=0.3$.
Positive Feature Contrastive Loss.
To encourage compact and discriminative class-wise clustering in the latent feature space, we introduce a contrastive loss. This loss enforces each positive feature $\mathbf{h}_{ic}$ to be closer to its own prototype than to others via a prototype-softmax:
@@
Total Objective.
The total loss can be expressed as:
@@
where $\lambda_{1},\lambda_{2}\geq 0$ balance the contributions of the regularizers.
FedNCAlign-ML
We propose FedNCAlign-ML, a framework that enhances the standard FL protocol with an NC-inspired local training strategy tailored for multi-label tasks and skewed local data distributions. In this framework, each client first encodes its input images into features, which are then disentangled into class-specific embeddings using LADM. These embeddings are aligned with a shared simplex ETF matrix that serves simultaneously as fixed classifier prototypes and as class-wise queries for LADM, thereby enforcing a consistent global geometry across clients despite label imbalance. To further improve representation quality, we incorporate two regularizers: (i) a negative rejection loss to suppress spurious alignments of irrelevant features, and (ii) a positive contrastive loss to encourage compact intra-class clustering and clear inter-class separation. The novelty of FedNCAlign-ML lies in this local training design, which integrates disentangled class-wise embeddings, ETF-based alignment, and geometry-aware regularization. The complete procedure is summarized in Algorithm 1.
Algorithm 1 FedNCAlign-ML: Federated Neural Collapse Alignment for Multi-Label Learning
@@
7:  $w_{t+1}\leftarrow\textsc{ModelAggregation}(\{w_{t+1}^{k}\}_{k=1}^{K})$
@@
12:  for $e=1$ to $E$ do $\triangleright$ local epochs
@@
15:     $\mathbf{H}\leftarrow\textsc{LADM}(\mathbf{f},\mathbf{M})$
@@
$\beta=0.5,\gamma=0.5$ $\beta=0.1,\gamma=0.5$
macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
Centralized (Upperbound) 91.17 61.40 91.04 62.22 91.17 61.40 91.04 62.22
FedAvg (McMahan et al.,, 2017) 82.46 40.65 81.83 41.24 78.84 31.15 77.00 34.98
FedCurv (Shoham et al.,, 2019) 82.19 39.36 81.81 39.61 79.03 32.31 77.97 36.13
FedProx (Li et al.,, 2020) 82.23 39.37 81.68 39.27 79.40 31.01 76.69 35.73
SCAFFOLD (Karimireddy et al.,, 2020) 82.01 38.41 81.49 38.83 79.29 32.38 77.74 35.90
SphereFed (Dong et al.,, 2022) 84.85 47.58 85.04 47.29 83.27 39.52 79.45 40.77
FedLGT (Liu et al.,, 2024) 84.24 45.45 84.16 45.99 81.44 35.95 81.49 38.78
@@
Table 1: Comparisons on the multi-label CIFAR-10 dataset (Krizhevsky et al.,, 2009). Non-IID data distributions across clients are simulated using the Dirichlet concentration parameter ($\beta$) and the class-missing ratio ($\gamma$).
$\beta=0.5,\gamma=0.71$ $\beta=0.1,\gamma=0.71$
macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
Centralized (Upperbound) 91.86 63.49 94.52 73.30 91.86 63.49 94.52 73.30
FedAvg (McMahan et al.,, 2017) 90.18 53.96 92.68 67.53 83.95 40.69 86.68 63.03
FedCurv (Shoham et al.,, 2019) 90.31 56.17 92.64 68.11 86.20 42.19 89.54 64.76
FedProx (Li et al.,, 2020) 90.16 55.85 92.44 67.95 83.73 42.98 87.96 63.50
SCAFFOLD (Karimireddy et al.,, 2020) 89.56 54.88 92.13 67.59 85.33 43.60 89.41 64.75
SphereFed (Dong et al.,, 2022) 85.36 44.25 89.46 53.40 83.37 41.90 88.92 50.94
FedLGT (Liu et al.,, 2024) 87.33 56.24 90.84 68.25 84.80 44.09 88.70 57.35
@@
Table 2: Comparisons on the multi-label DermaMNIST dataset (Yang et al., 2023a, ).
$\beta=0.05,\gamma=0.5$ $\beta=0.01,\gamma=0.5$
macro-AUC macro-F1 AUC F1 macro-AUC macro-F1 AUC F1
Centralized (Upperbound) 95.66 74.77 96.22 77.15 95.66 74.77 96.22 77.15
FedAvg (McMahan et al.,, 2017) 93.72 57.00 94.20 64.62 93.29 49.16 93.90 61.08
FedCurv (Shoham et al.,, 2019) 93.69 57.32 94.40 64.74 93.33 48.93 93.87 61.35
FedProx (Li et al.,, 2020) 93.70 57.08 94.26 64.46 93.05 48.99 93.74 61.16
SCAFFOLD (Karimireddy et al.,, 2020) 93.07 58.20 93.97 65.86 93.44 50.89 93.93 62.10
SphereFed (Dong et al.,, 2022) 85.72 33.71 87.38 38.19 86.85 32.44 87.94 36.18
FedLGT (Liu et al.,, 2024) 91.07 61.40 91.39 65.88 91.20 56.59 91.20 64.58
@@
Table 3: Comparison of related FL methods with ours on the multi-label PASCAL VOC 2012 dataset (Everingham et al.,, 2010).
Experiment
Dataset and Evaluation Metric
@@
Datasets: We conduct experiments on both general computer vision (CV) datasets and medical imaging datasets to evaluate the effectiveness and real-world applicability of the proposed method. For general CV tasks, we use CIFAR-10 (Krizhevsky et al.,, 2009) and PASCAL VOC (Everingham et al.,, 2010). Since CIFAR-10 is originally a multi-class dataset, we adopt the method used by Li et al., 2023a to transform it into a multi-label dataset. This is done by merging multiple images into a single composite image, with the combined labels serving as the new ground truth. For medical imaging tasks, we use DermaMNIST (Yang et al., 2023a, ) and ChestX-ray14 (Wang et al.,, 2017). DermaMNIST, which originally contains 7 skin disease categories in a multi-class format, is converted to a multi-label dataset using the same strategy applied to CIFAR-10. ChestX-ray14 is naturally a multi-label dataset, comprising 14 thoracic disease categories and an additional “No Finding” label. Since a significant portion of the dataset, approximately four-sevenths of the training data, are “No Finding” samples (negative cases with all-zero labels), we distribute these samples evenly across all clients. This setup mimics a realistic clinical scenario in which healthy cases are prevalent, while disease cases are relatively rare and unevenly distributed across clients. Detailed information about the datasets and local data distributions under various experimental settings is provided in the Appendix.
@@
Evaluation Metric: Given our focus on label-skewed data distributions, where certain classes may be missing or underrepresented, we report both overall (micro) and class-wise (macro) performance metrics. Following standard practices across different datasets, we report AUC and F1 scores for CIFAR-10, PASCAL VOC, and DermaMNIST, and AUC for ChestX-ray14. Together, these metrics offer a balanced and comprehensive evaluation by capturing both overall performance and class-specific behaviour.
Task Setup and Implementation Details
Implementation Details: For all experiments, we use ResNet-18 (He et al.,, 2016) as the feature extractor. Each FL global model is trained over 100 communication rounds, with one epoch of local training per round. The final model is selected based on its best performance on the validation set. We use a batch size of 32 and initialise the learning rate at $1\times 10^{-4}$, using the AdamW optimiser with a weight decay of 0.01. Models for the CIFAR-10 dataset are trained from scratch, while models for other datasets are initialised with ImageNet-pretrained weights. During training, the regularisation hyperparameter $\lambda_{1}$ is empirically set to either 1 or 0.01, depending on the background complexity of the dataset. For datasets with complex backgrounds, where disentangling class-specific features is more challenging, a smaller value ($\lambda_{1}=0.01$) is used to reduce potential interference with positive feature learning. The other regularisation hyperparameter, $\lambda_{2}$, is fixed at 1 across all experiments.
Performance Comparison
To evaluate the effectiveness of the proposed method, we compare it with other methods across four datasets, as shown in Tables 3, 3, 3, and 4. As observed in these tables, our method consistently achieves the best performance in most cases across different datasets. The improvement is particularly notable on naturally long-tailed datasets such as DermaMNIST and ChestX-ray14, where our method yields substantial gains in class-wise performance, indicating its ability to produce more balanced predictions. Specifically, under the $\beta=0.1$ setting on DermaMNIST, our method outperforms the second-best approach by 1.49% in class-wise AUC and 7.29% in class-wise F1 score. For PASCAL VOC, when each client is limited to a maximum of half the total classes, our method achieves improvements of 2.94% and 5.49% in class-wise F1 score under the $\beta=0.05$ and $\beta=0.01$ settings, respectively. For ChestX-ray14 under the $\beta=0.5$ and $\beta=0.1$ settings, our method improves class-wise AUC by 0.64% and 1.31%, respectively. Overall, across a range of non-IID configurations with limited class availability per client, our method consistently outperforms existing FL approaches, demonstrating strong robustness to various label-skewed data distributions.
@@
(a) FC classifier
(b) LADM + fixed ETF classifier
(c) LADM + ETF + Dual regularization
Figure 4: Pair-wise cosine similarity between test data class-wise average feature prototypes.
$\beta=0.5,\gamma=0.5$ $\beta=0.1,\gamma=0.5$
macro-AUC AUC macro-AUC AUC
Centralized 71.44 79.62 71.44 79.62
FedAvg 69.07 73.28 69.84 78.00
FedCurv 68.12 72.15 69.46 77.63
FedProx 69.23 72.45 69.23 77.69
SCAFFOLD 69.02 72.21 66.71 78.69
SphereFed 57.48 70.26 61.80 74.59
FedLGT 69.76 72.41 70.02 77.93

### v3

@@ 4.1 Label-Aware Disentanglement Module
$\mathbf{h}_{ic}\;=\;\operatorname{MultiHeadAttn}\bigl(\mathbf{m}_{c}^{\top},\,\tilde{\mathbf{Z}}_{i},\,\tilde{\mathbf{Z}}_{i}\bigr),\quad c\in\mathcal{C}.$ (7)
The resulting features are stacked to form the class-feature matrix:
@@
4.2 Neural Collapse-Inspired Feature Alignment
After disentangling image-level features into class-wise representations with LADM, each sample $i\in[N_{k}]$ yields a feature $\mathbf{h}_{ic}\in\mathbb{R}^{d}$ for every class $c\in\mathcal{C}$. To prevent client-specific classifier drift, we fix the classifier to the global simplex ETF, which also serves as the LADM query matrix. This shared classifier enhances class separability and mitigates model drift caused by class imbalance and missing labels. Given $\mathbf{h}_{ic}$ and its corresponding prototype $\mathbf{m}_{c}$, we compute the class logit and apply the sigmoid function to obtain a binary prediction:
@@ 4.2 Neural Collapse-Inspired Feature Alignment
and train with a binary cross-entropy (BCE) loss:
@@
Here, $y_{ic}\in\{0,1\}$ is the ground-truth multi-label indicator, and $\sigma(\cdot)$ is the sigmoid function.
4.3 Reducing Noise and Improving Clustering
Each sample produces $C$ class-wise features $\{\mathbf{h}_{ic}\}_{c=1}^{C}$. Let the positive and negative class sets for sample $i$ be defined as $\mathcal{C}_{i}^{+}=\{\,c\mid y_{ic}=1\,\}$ and $\mathcal{C}_{i}^{-}=\{\,c\mid y_{ic}=0\,\}$, respectively. We introduce two regularization terms: one to suppress noise from negative features, and another to promote compact clustering of positive features.
Negative Feature Rejection Loss. While $\mathcal{L}_{\text{BCE}}$ discourages alignment between a negative feature $\mathbf{h}_{ic}$ and its corresponding prototype $\mathbf{m}_{c}$ when $y_{ic}=0$, it does not prevent $\mathbf{h}_{ic}$ from spuriously aligning with other class prototypes. To address this, we introduce a penalty on high similarity between each negative feature and all non-self prototypes:
@@ 4.3 Reducing Noise and Improving Clustering
and define the negative feature rejection loss as
$\begin{split}\mathcal{L}_{\text{Neg}}&=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{-}|}\sum_{c\in\mathcal{C}_{i}^{-}}\\[4.0pt] &\quad\times\frac{1}{C-1}\sum_{\begin{subarray}{c}r=1\\ r\neq c\end{subarray}}^{C}\mathbb{I}\!\big(\hat{s}_{icr}>\tau\big)\,\log\!\big(1-\hat{s}_{icr}\big)\end{split}$ (12)
The indicator $\mathbb{I}(\cdot)$ filters out low-similarity pairs, ensuring that only confident negatives contribute to the loss. In our experiments, we set $\tau=0.3$.
Positive Feature Contrastive Loss. To encourage compact and discriminative class-wise clustering in the latent feature space, we introduce a contrastive loss. This loss drives each positive feature $\mathbf{h}_{ic}$ to be closer to its own prototype than to others through a prototype-based softmax:
@@
Total Objective. The total training loss is defined as:
@@
where $\lambda_{1},\lambda_{2}\geq 0$ are weighting coefficients that balance the contributions of the regularization terms.
4.4 FedNCA-ML
In summary, we propose FedNCA-ML, a framework that enhances the standard FL protocol with an NC-inspired local training strategy for multi-label tasks under skewed client data distributions. Each client encodes images into features that are disentangled into class-specific embeddings via LADM, which are then aligned with a shared simplex ETF matrix serving as both fixed classifier prototypes and class-wise queries. This alignment enforces a consistent global geometry across clients despite label imbalance. Two regularizers further refine representation quality: a negative rejection loss to suppress spurious alignments, and a positive contrastive loss to promote compact intra-class clustering and clear inter-class separation. This geometry-aware training framework unifies disentangled embeddings, ETF-based alignment, and structure-preserving regularization. The overall procedure is summarized in Algorithm 1.
Algorithm 1 FedNCA-ML
@@ 4.4 FedNCA-ML
7:  $w_{t+1}\leftarrow\frac{1}{K}\sum_{k=1}^{K}w_{t+1}^{k}$ $\triangleright$ model aggregation
@@
12:  for $e=0$ to E1E-1 do $\triangleright$ local epochs
@@
15:     $\mathbf{H}\leftarrow\textsc{LADM}(\mathbf{F},\mathbf{M})$ $\triangleright$ Eqs. 345678
@@
Table 1: Comparisons on multi-label CIFAR-10 (krizhevsky2009learning) with the class presence ratio ($\gamma$) set to 0.5 ($\leq$ 5 of 10 classes per client). Non-IID client distributions are simulated using the Dirichlet factor ($\beta$). Class-wise (macro) and instance-wise (micro) performance are reported.
$\beta=0.5$ $\beta=0.1$
macro-AUC macro-F1 micro-AUC micro-F1 macro-AUC macro-F1 micro-AUC micro-F1
Centralized (Upperbound) 92.20 $\pm$ 0.36 61.30 $\pm$ 0.94 90.04 $\pm$ 0.65 62.02 $\pm$ 0.62 92.20 $\pm$ 0.36 61.30 $\pm$ 0.94 90.04 $\pm$ 0.65 62.02 $\pm$ 0.62
FedAvg (mcmahan2017communication) 82.29 $\pm$ 0.46 39.47 $\pm$ 1.59 81.48 $\pm$ 0.78 40.26 $\pm$ 1.09 78.92 $\pm$ 0.39 31.17 $\pm$ 0.51 77.62 $\pm$ 0.24 35.07 $\pm$ 0.47
FedCurv (shoham2019overcoming) 82.53 $\pm$ 0.30 39.96 $\pm$ 1.28 82.10 $\pm$ 0.29 40.34 $\pm$ 1.04 79.06 $\pm$ 0.51 31.00 $\pm$ 1.21 77.46 $\pm$ 0.47 35.03 $\pm$ 0.54
FedProx (li2020federated) 82.45 $\pm$ 0.34 39.22 $\pm$ 0.74 81.77 $\pm$ 0.48 39.66 $\pm$ 0.53 78.82 $\pm$ 0.55 30.74 $\pm$ 0.47 77.40 $\pm$ 0.31 35.02 $\pm$ 0.27
SCAFFOLD (karimireddy2020scaffold) 82.51 $\pm$ 0.44 39.98 $\pm$ 1.41 82.08 $\pm$ 0.53 40.26 $\pm$ 1.26 79.00 $\pm$ 0.23 31.38 $\pm$ 0.31 77.72 $\pm$ 0.15 35.54 $\pm$ 0.16
SphereFed (dong2022spherefed) 83.63 $\pm$ 1.50 42.58 $\pm$ 3.20 83.50 $\pm$ 1.87 43.18 $\pm$ 2.34 80.62 $\pm$ 1.46 36.83 $\pm$ 1.39 78.37 $\pm$ 0.95 38.18 $\pm$ 1.40
FedLGT (liu2024language) 83.52 $\pm$ 0.68 43.60 $\pm$ 1.68 83.36 $\pm$ 0.71 44.03 $\pm$ 1.71 80.54 $\pm$ 0.43 36.30 $\pm$ 0.82 80.65 $\pm$ 0.68 39.24 $\pm$ 0.71
@@
Table 2: Comparisons on multi-label DermaMNIST (yang2023medmnist) with the class presence ratio ($\gamma$) set to 0.71 ($\leq$ 5 of 7 classes per client).
$\beta=0.5$ $\beta=0.1$
macro-AUC macro-F1 micro-AUC micro-F1 macro-AUC macro-F1 micro-AUC micro-F1
Centralized (Upperbound) 91.83 $\pm$ 0.40 64.38 $\pm$ 1.01 94.48 $\pm$ 0.25 74.12 $\pm$ 0.70 91.83 $\pm$ 0.40 64.38 $\pm$ 1.01 94.48 $\pm$ 0.25 74.12 $\pm$ 0.70
FedAvg (mcmahan2017communication) 89.72 $\pm$ 0.29 54.88 $\pm$ 0.97 92.51 $\pm$ 0.36 68.29 $\pm$ 0.45 83.88 $\pm$ 1.21 42.79 $\pm$ 2.03 87.23 $\pm$ 0.86 62.81 $\pm$ 0.83
FedCurv (shoham2019overcoming) 89.48 $\pm$ 0.18 54.92 $\pm$ 1.11 92.26 $\pm$ 0.36 67.97 $\pm$ 0.92 85.53 $\pm$ 1.37 43.45 $\pm$ 1.20 88.87 $\pm$ 0.50 64.37 $\pm$ 0.26
FedProx (li2020federated) 89.69 $\pm$ 0.31 55.78 $\pm$ 0.94 92.06 $\pm$ 0.46 68.01 $\pm$ 1.46 84.45 $\pm$ 1.34 43.54 $\pm$ 1.35 88.28 $\pm$ 1.07 63.25 $\pm$ 0.63
SCAFFOLD (karimireddy2020scaffold) 89.72 $\pm$ 0.49 55.85 $\pm$ 0.68 92.45 $\pm$ 0.32 68.09 $\pm$ 0.59 84.59 $\pm$ 1.03 43.20 $\pm$ 1.00 88.91 $\pm$ 0.90 63.82 $\pm$ 1.16
SphereFed (dong2022spherefed) 85.09 $\pm$ 0.87 43.41 $\pm$ 1.73 89.65 $\pm$ 0.55 65.24 $\pm$ 0.58 81.78 $\pm$ 1.14 40.90 $\pm$ 1.79 86.59 $\pm$ 1.87 54.94 $\pm$ 1.12
FedLGT (liu2024language) 87.63 $\pm$ 0.71 55.82 $\pm$ 1.34 91.19 $\pm$ 0.51 67.93 $\pm$ 0.45 84.91 $\pm$ 0.79 45.61 $\pm$ 1.16 89.42 $\pm$ 0.65 61.15 $\pm$ 2.43
@@
Table 3: Comparisons on PASCAL VOC (everingham2010pascal) with the class presence ratio ($\gamma$) set to 0.5 ($\leq$ 10 of 20 classes per client).
$\beta=0.05$ $\beta=0.01$
macro-AUC macro-F1 micro-AUC micro-F1 macro-AUC macro-F1 micro-AUC micro-F1
Centralized (Upperbound) 95.48 $\pm$ 0.11 74.61 $\pm$ 0.10 96.11 $\pm$ 0.12 76.94 $\pm$ 0.09 95.48 $\pm$ 0.11 74.61 $\pm$ 0.10 96.11 $\pm$ 0.12 76.94 $\pm$ 0.09
FedAvg (mcmahan2017communication) 93.54 $\pm$ 0.23 57.57 $\pm$ 0.76 94.13 $\pm$ 0.33 64.67 $\pm$ 0.35 93.31 $\pm$ 0.11 47.73 $\pm$ 1.03 93.05 $\pm$ 0.19 60.46 $\pm$ 0.43
FedCurv (shoham2019overcoming) 93.53 $\pm$ 0.33 57.36 $\pm$ 0.41 94.10 $\pm$ 0.20 64.60 $\pm$ 0.26 93.13 $\pm$ 0.29 48.54 $\pm$ 0.68 93.81 $\pm$ 0.09 60.49 $\pm$ 1.16
FedProx (li2020federated) 93.55 $\pm$ 0.20 57.04 $\pm$ 0.24 94.04 $\pm$ 0.17 64.43 $\pm$ 0.29 93.14 $\pm$ 0.25 49.49 $\pm$ 0.68 93.80 $\pm$ 0.17 62.18 $\pm$ 0.51
SCAFFOLD (karimireddy2020scaffold) 93.17 $\pm$ 0.20 57.44 $\pm$ 0.33 94.03 $\pm$ 0.11 64.95 $\pm$ 0.21 93.41 $\pm$ 0.17 49.64 $\pm$ 0.77 93.60 $\pm$ 0.21 61.44 $\pm$ 0.56
SphereFed (dong2022spherefed) 83.72 $\pm$ 1.82 33.49 $\pm$ 1.15 84.38 $\pm$ 1.10 38.19 $\pm$ 1.17 84.25 $\pm$ 2.44 32.44 $\pm$ 0.21 85.94 $\pm$ 3.67 35.18 $\pm$ 1.30
FedLGT (liu2024language) 91.93 $\pm$ 0.42 62.11 $\pm$ 0.44 91.75 $\pm$ 0.42 67.58 $\pm$ 0.45 91.25 $\pm$ 0.27 56.53 $\pm$ 2.42 91.46 $\pm$ 0.41 63.51 $\pm$ 1.73
@@
5 Experiments
5.1 Dataset and Evaluation Metric
@@
Datasets: We conduct experiments on both general computer vision (CV) datasets and medical imaging datasets to evaluate the effectiveness and real-world applicability of the proposed method. For general CV tasks, we use CIFAR-10 (krizhevsky2009learning) and PASCAL VOC (everingham2010pascal). Since CIFAR-10 is originally a multi-class dataset, we adopt the method used by li2023neural to transform it into a multi-label dataset. This is done by merging multiple images into a single composite image, with the combined labels serving as the new ground truth. For medical imaging tasks, we use DermaMNIST (yang2023medmnist) and ChestX-ray14 (wang2017chestxray). DermaMNIST, which originally contains 7 skin disease categories in a multi-class format, is converted to a multi-label dataset using the same strategy applied to CIFAR-10. ChestX-ray14 is naturally a multi-label dataset, comprising 14 thoracic disease categories and an additional “No Finding” label. Since a significant portion of the dataset, 57% of the training data, is “No Finding” samples (negative cases with all-zero labels), we distribute these samples evenly across all clients. This setup mimics a realistic clinical scenario in which healthy cases are prevalent, while disease cases are relatively rare and unevenly distributed across clients. Detailed information about the datasets and local data distributions under various experimental settings is provided in the Appendix C.
@@
Evaluation Metric: Given our focus on label-skewed data distributions, we report both instance-wise (micro) and class-wise (macro) performance metrics. The macro metric provides a balanced evaluation across classes, mitigating bias toward frequent categories. Following standard practice, we report AUC and F1 scores for CIFAR-10, PASCAL VOC, and DermaMNIST, and AUC for ChestX-ray14. Together, these metrics comprehensively evaluate overall performance and class-level behaviour.
@@
5.2 Task Setup and Implementation Details
Implementation Details: All experiments adopt ResNet-18 (he2016deep) as the feature extractor. Each FL global model is trained for 100 communication rounds, with one local epoch per round, and the final model is selected based on its best validation performance. We use a batch size of 32, an initial learning rate of $1\times 10^{-4}$, and the AdamW optimizer with a weight decay of 0.01. Models on CIFAR-10 are trained from scratch, while models on other datasets are initialised with ImageNet-pretrained weights. The negative feature rejection regularization coefficient, $\lambda_{1}$, is empirically set to 1 for CIFAR-10 and 0.01 for other datasets. The positive feature contrastive loss coefficient, $\lambda_{2}$, is set to 1 across all experiments. To ensure fair comparison, all experiments are repeated three times with different random seeds, and the mean and standard deviation are reported.
5.3 Performance Comparison
To evaluate the effectiveness of the proposed method, we compare FedNCA-ML with state-of-the-art approaches across four datasets under eight label-skewed FL settings. The results, summarized in Tables 3, 3, 3, and 4, show that FedNCA-ML consistently achieves the best class-wise performance in most cases, demonstrating its strong capability for balanced and generalized predictions under heterogeneous FL conditions. Specifically, on multi-label CIFAR-10, under non-IID Dirichlet settings of $\beta=0.5$ ($\beta=0.1$) with a maximum of 5 out of 10 classes per client, FedNCA-ML surpasses the second-best approach by 3.92% (3.18%) in class-wise AUC and 4.57% (1.26%) in class-wise F1 score. On multi-label DermaMNIST, under $\beta=0.5$ ($\beta=0.1$) with up to 5 of 7 classes per client, it achieves improvements of 0.40% (0.46%) in AUC and 0.77% (4.93%) in F1 score. For PASCAL VOC, under more challenging non-IID settings of $\beta=0.05$ ($\beta=0.01$) with up to 10 of 20 classes per client, our method outperforms the second-best approach by 2.17% (4.55%) in class-wise F1 score.
On ChestX-ray14, under non-IID Dirichlet settings of $\beta=0.5$ and $\beta=0.1$ with up to 7 of 14 disease classes per client, FedNCA-ML improves class-wise AUC by 0.69% and 1.21%, respectively. However, our method does not outperform others in overall AUC. ChestX-ray14 is highly imbalanced with 57% of training and 38% of testing samples labeled as “No Finding”. This severe imbalance leads many methods to overpredict the majority class, as reflected in a large gap between class-wise and overall AUC, indicating bias toward majority classes and degraded performance on minority (disease) classes. In medical diagnosis, false positives are generally more tolerable than false negatives, especially for rare diseases. The superior class-wise AUC and smaller gap between overall and class-wise AUC achieved by FedNCA-ML suggest more balanced predictions and better recognition of minority disease classes.
Table 4: Comparisons on ChestX-ray14 (wang2017chestxray) with $\gamma$ set to 0.5.
$\beta=0.5$ $\beta=0.1$
macro-AUC micro-AUC macro-AUC micro-AUC
Centralized 71.66 $\pm$ 0.12 79.54 $\pm$ 0.36 71.66 $\pm$ 0.12 79.54 $\pm$ 0.36
FedAvg 69.02 $\pm$ 0.24 73.18 $\pm$ 0.15 69.05 $\pm$ 0.78 77.90 $\pm$ 0.20
FedCurv 68.72 $\pm$ 1.19 72.15 $\pm$ 0.17 69.34 $\pm$ 0.48 77.36 $\pm$ 0.27
FedProx 69.02 $\pm$ 0.21 72.04 $\pm$ 0.59 69.59 $\pm$ 0.23 77.43 $\pm$ 0.24
SCAFFOLD 69.42 $\pm$ 0.39 72.49 $\pm$ 0.38 67.45 $\pm$ 0.54 77.90 $\pm$ 0.69
SphereFed 58.35 $\pm$ 0.57 69.51 $\pm$ 0.56 61.96 $\pm$ 0.16 73.59 $\pm$ 1.02
FedLGT 69.86 $\pm$ 0.76 72.27 $\pm$ 1.07 70.16 $\pm$ 0.37 77.67 $\pm$ 0.57